OpenStudy (anonymous):

MEDAL & FAN!! 6. The length of a rectangle is 5 mm longer than its width. Its perimeter is more than 34mm. Let w equal the width of the rectangle. a. Write an expression for the length in terms of the width. b. Use expressions for the length and width to write an inequality for the perimeter, on the basis of the given information. c. Solve the inequality, clearly indicating the width of the rectangle.

OpenStudy (anonymous):

@stamp do you think you can help with this one? :/

OpenStudy (anonymous):

@horsegirl27 can you help?

OpenStudy (horsegirl27):

P>34 34>w+5 is part A

OpenStudy (anonymous):

Is there an image

OpenStudy (horsegirl27):

I think

OpenStudy (stamp):

6. The length of a rectangle is 5 mm longer than its width. Its perimeter is more than 34mm. Let w equal the width of the rectangle. Break down the sentences. The length L of a rectangle is 5 mm longer than its width W. OR L = W + 5 It's perimeter P is more than ( > ) 34 mm. OR P > 34

OpenStudy (stamp):

If we had a way to relate the variables P, L, and W, we can look at the problem through a different lens. Fortunately, we know a relation. Perimeter = L + L + W + W, OR P = 2L + 2W OR P = 2(L + W)

OpenStudy (stamp):

a. Write an expression for the length in terms of the width. We did this earlier by saying L = W + 5

OpenStudy (stamp):

b. Use expressions for the length and width to write an inequality for the perimeter, on the basis of the given information. We said earlier that P > 34 and P = 2(L + W). Let us combine the two thoughts to declare 2(L + W) > 34

OpenStudy (stamp):

c. Solve the inequality, clearly indicating the width of the rectangle. Well, we know two things now. L = W + 5 and 2(L + W) > 34. Again we combine the two thoughts by substituting L in the second statement to realize 2(W + 5 + W) > 34

OpenStudy (stamp):

You can solve for W by yourself